OpenStudy (anonymous):

An astronaut who weighs 803.9 N on the sur- face of the earth lifts off from planet Zuton in a space ship. The free-fall acceleration on Zuton is 6.3 m/s2. At the moment of lift off the space ship experiences an acceleration of 4.8 m/s2. The acceleration of gravity is 9.8 m/s2 . What is the magnitude of the force that the space ship exerts on the astronaut? Answer in units of N

OpenStudy (anonymous):

Weight on earth = 803.9 N Weight = m*g 803.9 =m*g 803.9 =m*9.8 803.9/9.8 = m mass = 82 kg  If the space ship is moving upwards, then the upwards force must be larger than the downwards force in order to lift off. In other words, the Force of liftoff must be greater than the force of gravity on the planet Zuton. Sum of the forces in the Y-direction = ma Sum F =ma F(liftoff) - F(gravity of Zulton) = (mass)(acceleration of the total system) F(liftoff) - (Mass)(Gravity of Zulton) = (mass)(acceleration of the total system) F(liftoff) - (82kg)(6.3m/s^2) = (82kg)(4.8m/s^2) F(liftoff) - 516.6N = 393.6N F(liftoff) = 393.6N + 516.6N F(liftoff) = 910N Hope it's right :)